OpenStudy (jazzyc15):

f(x) = 7x − 13. Find f−1(x).

OpenStudy (mathmate):

hint: two steps 1. exchange x and y 2. solve for y in terms of x. Example: y=g(x)=3x-2 1. exchange x and y x=3y-2 3y=x+2 y=(x+2)/3 \(\Large g^{-1}(x)=\frac{x+2}{3}\)